# On Discounted Infinite-Time Mean Field Games

### v3

@@ 2 Preliminaries
$\begin{cases}\mathrm{d}X_{t}=B(t,X_{t},Y_{t},\mathcal{L}_{X_{t}})dt+\mathrm{d}B_{t},\\ \mathrm{d}Y_{t}=-F(t,X_{t},Y_{t},\mathcal{L}_{X_{t}})dt+Z_{t}\mathrm{d}B_{t},\\ X_{0}=\xi.\end{cases}$ (2.3)
@@
$|B(t,x,y,\mu)$ $-B(t,x^{\prime},y^{\prime},\mu^{\prime})|+|F(t,x,y,\mu)-F(t,x^{\prime},y^{\prime},\mu^{\prime})|$ (2.4)
@@
Finally, we consider the space $\Theta\triangleq[0,+\infty)\times\mathbb{R}\times{\cal P}_{2}$, and let ${\cal C}^{1,2,1}(\Theta)$ denote the set of continuous functions $U:\Theta\to\mathbb{R}$ which has the following continuous derivatives: $\partial_{t}U$, $\partial_{x}U$, $\partial_{xx}U$, $\partial_{\mu}U$, $\partial_{x\mu}U$, $\partial_{\tilde{x}\mu}U.$ One crucial property of functions $U\in{\cal C}^{1,2,1}(\Theta)$ is the Itô’s formula[2, 5]. For $i=1,2$, let $dX^{i}_{t}\triangleq b^{i}_{t}dt+\sigma^{i}_{t}dB_{t},$ where $b^{i}:[0,+\infty)\times\Omega\to\mathbb{R}$ and $\sigma^{i}:[0,+\infty)\times\Omega\to\mathbb{R}$ are $\mathbb{F}$-progressively measurable and bounded (for simplicity), and $\rho_{t}:={\cal L}_{X^{2}_{t}}$. Fix $T>0$ and let all conditions be restricted to the interval [0,T]. Then we have
$\begin{split}\mathrm{d}U(t,X^{1}_{t},\rho_{t})=&\Big{[}\partial_{t}U+\partial_{x}U\cdot b^{1}_{t}+\frac{1}{2}\hbox{\rm tr$\,$}\big{(}\partial_{xx}U[\sigma_{t}^{1}]^{2}\big{)}\Big{]}(t,X^{1}_{t},\rho_{t})\mathrm{d}t\\ &+\Big{(}\tilde{\mathbb{E}}_{{\cal F}_{t}}\big{[}\partial_{\mu}U(t,X^{1}_{t},\rho_{t},\tilde{X}^{2}_{t})(\tilde{b}^{2}_{t})+\frac{1}{2}\partial_{\tilde{x}}\partial_{\mu}U(t,X^{1}_{t},\rho_{t},\tilde{X}^{2}_{t})[\tilde{\sigma}_{t}^{2}]^{2}\big{]}\Big{)}\mathrm{d}t\\ &+\partial_{x}U(t,X^{1}_{t},\rho_{t})\sigma_{t}^{1}\mathrm{d}B_{t}.\end{split}$ (2.7)
@@ 5 Master equation representation
Since $b_{1}+|b_{2}|<\frac{r}{2}$, if we further require $a_{1}>0,a_{2}>0$, the SDE (5.3) admits a unique solution in $L_{r}^{2}$.
@@ 6.1 Yamada-Watanabe theorem for infinite-time FBSDEs
$\begin{cases}\mathrm{d}X_{t}=B(t,X_{t},Y_{t},\mathcal{L}_{X_{t}})dt+\mathrm{d}B_{t},\\ \mathrm{d}Y_{t}=-F(t,X_{t},Y_{t},\mathcal{L}_{X_{t}})dt+Z_{t}\mathrm{d}B_{t}.\end{cases}$ (6.4)
We know that under Assumption 2.1, the above FBSDE has unique solution $(X_{t},Y_{t},Z_{t})$ in $L_{r}^{2}$ with initial $X_{0}$. We then give the definitions of strong and weak uniqueness.
@@
Let us consider two filtered probability spaces $(\Omega^{i},\mathcal{F}^{i},\mathbb{P}^{i},\mathbb{F}^{i}),i=1,2$. Denote by $Q^{1}$ and $Q^{2}$ the distribution of $(X_{0}^{1},B_{t}^{1},X_{t}^{1},Y_{t}^{1},\int_{0}^{t}Z^{1}_{s}\mathrm{d}s)_{t\geq 0}$ and $(X_{0}^{2},B_{t}^{2},X_{t}^{2},Y_{t}^{2},\int_{0}^{t}Z^{2}_{s}\mathrm{d}s)_{t\geq 0}$ on $\Omega_{\text{canon}}=\Omega_{\text{input}}\times\Omega_{\text{output}}$, by $Q_{\text{input}}$ the common distribution of the processes $(X_{0}^{1},B_{t}^{1}),(X_{0}^{2},B_{t}^{2})$ on $\Omega_{\text{input}}$.
@@
$\begin{split}\mathbb{E}^{Q}\int_{0}^{\infty}\mathrm{e}^{rt}|z^{i}_{t}|\mathrm{d}t&\leq\mathop{\underline{\rm lim}}_{n\rightarrow\infty}\mathbb{E}^{Q}\int_{0}^{\infty}\mathrm{e}^{rt}|n(\zeta^{i}_{t}-\zeta^{i}_{(t-\frac{1}{n})+})|^{2}\mathrm{d}t\\ &=\mathop{\underline{\rm lim}}_{n\rightarrow\infty}n^{2}\mathbb{E}^{\mathbb{P}^{i}}\int_{0}^{\infty}\mathrm{e}^{rt}|\int_{(t-\frac{1}{n})+}^{t}Z_{s}^{i}\mathrm{d}s|^{2}\mathrm{d}t\\ &\leq\mathop{\underline{\rm lim}}_{n\rightarrow\infty}n\mathbb{E}^{\mathbb{P}^{i}}\int_{0}^{\infty}\mathrm{e}^{rt}\int_{(t-\frac{1}{n})+}^{t}|Z_{s}^{i}|^{2}\mathrm{d}s\mathrm{d}t\\ &\leq\mathbb{E}^{Q^{i}}\int_{0}^{\infty}\mathrm{e}^{rs}|Z_{s}^{i}|^{2}\mathrm{d}s,\end{split}$ (6.15)
@@
is measurable with respect to the completion of the $\sigma$-field ${\cal B}(\mathbb{R})\otimes{\cal K}_{t}$ under the probability measure $Q_{\text{input}}$, denoted $\overline{{\cal B}(\mathbb{R})\otimes{\cal K}_{t}}$. Indeed, let us consider $F\in{\cal B}(\mathbb{R})$, $G_{1}\in{\cal K}_{t}$ and $G_{2}\in\{w_{s}-w_{t};s\geq t\}$. Then, $\forall i\in\{1,2\}$
@@
Now we prove that $(w_{t})_{t\geq 0}$ is a $\{{\cal G}_{t}\}_{t\geq 0}$ Brownian motion. Let us consider $(A,A^{\prime})\in({\cal K}^{\prime}_{t})^{2},F\in{\cal B}(\mathbb{R}),G_{1}\in{\cal K}_{t}$ and $G_{2}\in\{w_{s}-w_{t};s\geq t\}$. Then,
@@
$\begin{cases}x^{i}_{t}=\xi+\int_{0}^{t}B(s,x^{i}_{s},y^{i}_{s},\mathcal{L}_{x^{i}_{s}})ds+w_{t},\\ y^{i}_{t}=y^{i}_{T}+\int_{t}^{T}F(s,x^{i}_{s},y^{i}_{s},\mathcal{L}_{x^{i}_{s}})ds-\int_{t}^{T}z^{i}_{s}\mathrm{d}w_{s},\\ \mathbb{E}^{Q}\left[\int_{0}^{\infty}\mathrm{e}^{-rt}\left(|x^{i}_{t}|^{2}+|y^{i}_{t}|^{2}+|z^{i}_{t}|^{2}\right)\mathrm{d}t\right]<\infty.\end{cases}$ (6.26)
@@ 6.3 Connection with distribution-dependent elliptic PDE
$(\mathcal{L}\mathcal{U})(x,\mu,\mathcal{U}(x,\mu))+F(x,\mu,\mathcal{U}(x,\mu))=0,$ (6.36)
@@
$\begin{split}(\mathcal{L}\Phi)(x,\mu,\Psi)\triangleq&\partial_{y}H(x,\mu,\Psi)\cdot\partial_{x}\Phi(x,\mu)+\frac{1}{2}\partial_{xx}\Phi(x,\mu)\\ &+\tilde{\mathbb{E}}\left[\frac{1}{2}\partial_{\tilde{x}}\partial_{\mu}\Phi(x,\mu,\tilde{\xi})+\partial_{\mu}\Phi(x,\mu,\tilde{\xi})\partial_{y}H(\tilde{\xi},\mu,\Psi(\tilde{\xi},\mu))\right],\end{split}$ (6.37)
@@
Let $\mathcal{U}\in C(\mathbb{R}\times\mathcal{P}_{2})$. Then $\mathcal{U}$ is called a viscosity subsolution (resp. supersolution) of PDE (6.36) if, whenever $\Psi\in C^{2,1}(\mathbb{R}\times\mathcal{P}_{2})$, and $(x^{0},\mu^{0})\in\mathbb{R}\times\mathcal{P}_{2}$ is a local maximum (resp. minimum) of $\mathcal{U}-\Psi$, we have
$(\mathcal{L}\Psi)(x^{0},\mu^{0},\mathcal{U}(x^{0},\mu^{0}))+F(x^{0},\mu^{0},\mathcal{U}(x^{0},\mu^{0}))\geq 0,$ (6.39)
@@
$(\mathcal{L}\Psi)(x^{0},\mu^{0},\mathcal{U}(x^{0},\mu^{0}))+F(x^{0},\mu^{0},\mathcal{U}(x^{0},\mu^{0}))\leq 0,$ (6.40)
). The function $\mathcal{U}$ is called a viscosity solution of PDE (6.36) if it is both a viscosity subsolution and a viscosity supersolution.
@@
Due to the uniqueness for solutions of FBSDE, it is not hard to see that for any $t\geq 0$, $\mathcal{V}(X_{t}^{x,\xi},\mathcal{L}_{X_{t}^{\xi}})=Y_{t}^{x,\xi}$. Let $\Psi\in C^{2,1}(\mathbb{R}\times\mathcal{P}_{2})$, and $(x^{0},\mu^{0})\in\mathbb{R}\times\mathcal{P}_{2}$ is a local maximum of $\mathcal{U}-\Psi$. We assume without loss of generality that $\mathcal{V}(x^{0},\mu^{0})=\Psi(x^{0},\mu^{0})$. We suppose that
$(\mathcal{L}\Psi)(x^{0},\mu^{0},\mathcal{V}(x^{0},\mu^{0}))+F(x^{0},\mu^{0},\mathcal{V}(x^{0},\mu^{0}))<0,$ (6.41)
@@
$\begin{cases}\mathcal{V}(x,\mu)\leq\Psi(x,\mu),\\ (\mathcal{L}\Psi)(x,\mu,\mathcal{V}(x,\mu))+F(x,\mu,\mathcal{V}(x,\mu))<0.\end{cases}$ (6.42)
Taking a initial state $\xi^{0}\in\mathbb{L}^{2}(\mathcal{F}_{0})$, we consider the processes $(X_{t}^{\xi^{0}},Y_{t}^{\xi^{0}},Z_{t}^{\xi^{0}})$ and $(X_{t}^{x^{0},\xi^{0}},Y_{t}^{x^{0},\xi^{0}},Z_{t}^{x^{0},\xi^{0}})$ which are solutions to FBSDEs (6.1) and (6.2). For some $T>0$, let $\tau$ denote the stopping time
$\tau\triangleq\inf\{t>0|(X_{t}^{x^{0},\xi^{0}},\mathcal{L}_{X_{t}^{\xi^{0}}})\in O\}\land T.$ (6.43)
We first note that the pair of processes
@@
$\begin{split}\bar{Y}_{t}=&\mathcal{V}(X^{x^{0},\xi^{0}}_{\tau},\mathcal{L}_{X^{\xi^{0}}_{\tau}})+\int_{t}^{T}I_{[0,\tau]}(s)F(X^{x^{0},\xi^{0}}_{s},\mathcal{L}_{X^{\xi^{0}}_{s}},\mathcal{V}(X^{x^{0},\xi^{0}}_{s},\mathcal{L}_{X^{\xi^{0}}_{s}}))\mathrm{d}s\\ &-\int_{t}^{T}\bar{Z}_{s}\mathrm{d}B_{s},\quad\quad 0\leq t\leq T.\end{split}$ (6.45)
@@
$(\hat{Y}_{t},\hat{Z}_{t})\triangleq(\Psi(X^{x^{0},\xi^{0}}_{t\wedge\tau},\mathcal{L}_{X^{\xi^{0}}_{t\wedge\tau}}),I_{\left[0,\tau\right]}(t)\partial_{x}\Psi(X^{x^{0},\xi^{0}}_{t\wedge\tau},\mathcal{L}_{X^{\xi^{0}}_{t\wedge\tau}}))$ (6.46)
@@
$\begin{split}\hat{Y}_{t}=&\Psi(X^{x^{0},\xi^{0}}_{\tau},\mathcal{L}_{X^{\xi^{0}}_{\tau}})-\int_{t}^{T}I_{[0,\tau]}(s)\mathcal{L}\Psi(X^{x^{0},\xi^{0}}_{s},\mathcal{L}_{X^{\xi^{0}}_{s}},\mathcal{V}(X^{x^{0},\xi^{0}}_{s},\mathcal{L}_{X^{\xi^{0}}_{s}}))\mathrm{d}s\\ &-\int_{t}^{T}\hat{Z}_{s}\mathrm{d}B_{s},\quad\quad 0\leq t\leq T.\end{split}$ (6.47)
@@
$\beta_{s}=-\mathcal{L}\Psi(X^{x^{0},\xi^{0}}_{s},\mathcal{L}_{X^{\xi^{0}}_{s}},\mathcal{V}(X^{x^{0},\xi^{0}}_{s},\mathcal{L}_{X^{\xi^{0}}_{s}}))-F(X^{x^{0},\xi^{0}}_{s},\mathcal{L}_{X^{\xi^{0}}_{s}},\mathcal{V}(X^{x^{0},\xi^{0}}_{s},\mathcal{L}_{X^{\xi^{0}}_{s}}))$ (6.48)
@@
$\begin{split}\tilde{Y}_{t}=&\Psi(X^{x^{0},\xi^{0}}_{\tau},\mathcal{L}_{X^{\xi^{0}}_{\tau}})-\mathcal{V}(X^{x^{0},\xi^{0}}_{\tau},\mathcal{L}_{X^{\xi^{0}}_{\tau}})+\int_{t}^{T}I_{[0,\tau]}(s)\beta_{s}\mathrm{d}s\\ &-\int_{t}^{T}\tilde{Z}_{s}\mathrm{d}B_{s},\quad\quad 0\leq t\leq T.\end{split}$ (6.50)
@@
$\tilde{Y}_{\tau}=\Psi(X^{x^{0},\xi^{0}}_{\tau},\mathcal{L}_{X^{\xi^{0}}_{\tau}})-\mathcal{V}(X^{x^{0},\xi^{0}}_{\tau},\mathcal{L}_{X^{\xi^{0}}_{\tau}})\geq 0,\quad\beta_{s}>0,\quad s\in[0,\tau].$ (6.52)

### v4

@@ 1 Introduction
Alternatively, if we define $\mathcal{V}(x,\mu)\triangleq Y^{x,\xi}_{0}$, we demonstrate that $\mathcal{V}$ serves as a viscosity solution to the distribution-dependent PDE below:
@@ 2 Preliminaries
$\begin{cases}\mathrm{d}X_{t}=G(t,X_{t},Y_{t},\mathcal{L}_{X_{t}})dt+\mathrm{d}B_{t},\\ \mathrm{d}Y_{t}=-F(t,X_{t},Y_{t},\mathcal{L}_{X_{t}})dt+Z_{t}\mathrm{d}B_{t},\\ X_{0}=\xi.\end{cases}$ (2.3)
@@
$|G(t,x,y,\mu)$ $-G(t,x^{\prime},y^{\prime},\mu^{\prime})|+|F(t,x,y,\mu)-F(t,x^{\prime},y^{\prime},\mu^{\prime})|$ (2.4)
@@
Finally, we consider the space $\Theta\triangleq[0,+\infty)\times\mathbb{R}\times{\cal P}_{2}$, and let ${\cal C}^{1,2,1}(\Theta)$ denote the set of continuous functions $U:\Theta\to\mathbb{R}$ which has the following continuous derivatives: $\partial_{t}U$, $\partial_{x}U$, $\partial_{xx}U$, $\partial_{\mu}U$, $\partial_{x\mu}U$, $\partial_{\tilde{x}\mu}U.$ One crucial property of functions $U\in{\cal C}^{1,2,1}(\Theta)$ is the Itô’s formula[2, 5]. For $i=1,2$, let $\mathrm{d}X^{i}_{t}\triangleq b^{i}_{t}\mathrm{d}t+\sigma^{i}_{t}\mathrm{d}B_{t},$ where $b^{i}:[0,+\infty)\times\Omega\to\mathbb{R}$ and $\sigma^{i}:[0,+\infty)\times\Omega\to\mathbb{R}$ are $\mathbb{F}$-progressively measurable and bounded (for simplicity), and $\rho_{t}\triangleq{\cal L}_{X^{2}_{t}}$. Fix $T>0$ and let all conditions be restricted to the interval $[0,T]$. Then we have
$\begin{split}\mathrm{d}U(t,X^{1}_{t},\rho_{t})=&\Big[\partial_{t}U+\partial_{x}U\cdot b^{1}_{t}+\frac{1}{2}\partial_{xx}U(\sigma_{t}^{1})^{2}\Big](t,X^{1}_{t},\rho_{t})\mathrm{d}t\\ &+\Big(\tilde{\mathbb{E}}_{{\cal F}_{t}}\big[\partial_{\mu}U(t,X^{1}_{t},\rho_{t},\tilde{X}^{2}_{t})(\tilde{b}^{2}_{t})+\frac{1}{2}\partial_{\tilde{x}}\partial_{\mu}U(t,X^{1}_{t},\rho_{t},\tilde{X}^{2}_{t})(\tilde{\sigma}_{t}^{2})^{2}\big]\Big)\mathrm{d}t\\ &+\partial_{x}U(t,X^{1}_{t},\rho_{t})\sigma_{t}^{1}\mathrm{d}B_{t}.\end{split}$ (2.7)
@@ 5 Master equation representation
Since the SDE (5.3) now becomes linear, if we further require $a_{1}>0,a_{2}>0$, it’s not hard to see that the SDE admits a unique solution in $L_{r}^{2}$.
@@ 6.1 Yamada-Watanabe theorem for infinite-time FBSDEs
$\begin{cases}\mathrm{d}X_{t}=G(t,X_{t},Y_{t},\mathcal{L}_{X_{t}})dt+\mathrm{d}B_{t},\\ \mathrm{d}Y_{t}=-F(t,X_{t},Y_{t},\mathcal{L}_{X_{t}})dt+Z_{t}\mathrm{d}B_{t}.\end{cases}$ (6.4)
We know that under Assumption 2.1, the above FBSDE has unique solution $(X_{t},Y_{t},Z_{t})$ in $L_{r}^{2}$ with initial $X_{0}\in\mathbb{L}^{2}({\cal F}_{0})$. We then give the definitions of strong and weak uniqueness.
@@
Let us consider two filtered probability spaces $(\Omega^{i},\mathcal{F}^{i},\mathbb{P}^{i},\mathbb{F}^{i})$ with identically distributed inputs $(X_{0}^{i},B^{i})$, $i=1,2$, on each of which a solution $(X_{t}^{i},Y_{t}^{i},\int_{0}^{t}Z^{i}_{s}\mathrm{d}s)_{t\geq 0}$ to the FBSDE (6.4) is defined. Denote by $Q^{1}$ and $Q^{2}$ the distribution of $(X_{0}^{1},B_{t}^{1},X_{t}^{1},Y_{t}^{1},\int_{0}^{t}Z^{1}_{s}\mathrm{d}s)_{t\geq 0}$ and $(X_{0}^{2},B_{t}^{2},X_{t}^{2},Y_{t}^{2},\int_{0}^{t}Z^{2}_{s}\mathrm{d}s)_{t\geq 0}$ on $\Omega_{\text{canon}}=\Omega_{\text{input}}\times\Omega_{\text{output}}$, by $Q_{\text{input}}$ the common distribution of the processes $(X_{0}^{1},B_{t}^{1}),(X_{0}^{2},B_{t}^{2})$ on $\Omega_{\text{input}}$.
@@
$\begin{split}\mathbb{E}^{Q}\int_{0}^{\infty}\mathrm{e}^{-rt}|z^{i}_{t}|\mathrm{d}t&\leq\mathop{\underline{\rm lim}}_{n\rightarrow\infty}\mathbb{E}^{Q}\int_{0}^{\infty}\mathrm{e}^{-rt}|n(\zeta^{i}_{t}-\zeta^{i}_{(t-\frac{1}{n})+})|^{2}\mathrm{d}t\\ &=\mathop{\underline{\rm lim}}_{n\rightarrow\infty}n^{2}\mathbb{E}^{\mathbb{P}^{i}}\int_{0}^{\infty}\mathrm{e}^{-rt}|\int_{(t-\frac{1}{n})+}^{t}Z_{s}^{i}\mathrm{d}s|^{2}\mathrm{d}t\\ &\leq\mathop{\underline{\rm lim}}_{n\rightarrow\infty}n\mathbb{E}^{\mathbb{P}^{i}}\int_{0}^{\infty}\mathrm{e}^{-rt}\int_{(t-\frac{1}{n})+}^{t}|Z_{s}^{i}|^{2}\mathrm{d}s\mathrm{d}t\\ &\leq\mathbb{E}^{Q^{i}}\int_{0}^{\infty}\mathrm{e}^{-rs}|Z_{s}^{i}|^{2}\mathrm{d}s,\end{split}$ (6.15)
@@
is measurable with respect to the completion of the $\sigma$-field ${\cal B}(\mathbb{R})\otimes{\cal K}_{t}$ under the probability measure $Q_{\text{input}}$, denoted $\overline{{\cal B}(\mathbb{R})\otimes{\cal K}_{t}}$. Indeed, let us consider $F\in{\cal B}(\mathbb{R})$, $G_{1}\in{\cal K}_{t}$ and $G_{2}\in\sigma\{w_{s}-w_{t};s\geq t\}$. Then, $\forall i\in\{1,2\}$
@@
Now we prove that $(w_{t})_{t\geq 0}$ is a $\{{\cal G}_{t}\}_{t\geq 0}$ Brownian motion. Let us consider $(A,A^{\prime})\in({\cal K}^{\prime}_{t})^{2},F\in{\cal B}(\mathbb{R}),G_{1}\in{\cal K}_{t}$ and $G_{2}\in\sigma\{w_{s}-w_{t};s\geq t\}$. Then,
@@
$\begin{cases}x^{i}_{t}=\xi+\int_{0}^{t}G(s,x^{i}_{s},y^{i}_{s},\mathcal{L}_{x^{i}_{s}})ds+w_{t},\\ y^{i}_{t}=y^{i}_{T}+\int_{t}^{T}F(s,x^{i}_{s},y^{i}_{s},\mathcal{L}_{x^{i}_{s}})ds-\int_{t}^{T}z^{i}_{s}\mathrm{d}w_{s},\\ \mathbb{E}^{Q}\left[\int_{0}^{\infty}\mathrm{e}^{-rt}\left(|x^{i}_{t}|^{2}+|y^{i}_{t}|^{2}+|z^{i}_{t}|^{2}\right)\mathrm{d}t\right]<\infty.\end{cases}$ (6.26)
@@ 6.3 Connection with distribution-dependent elliptic PDE
$(\mathcal{L}\mathcal{U})[\mathcal{U}](x,\mu)+F(x,\mu,\mathcal{U}(x,\mu))=0,$ (6.36)
@@
$\begin{split}(\mathcal{L}\Phi)[\Psi](x,\mu)\triangleq&\partial_{y}H(x,\mu,\Psi)\cdot\partial_{x}\Phi(x,\mu)+\frac{1}{2}\partial_{xx}\Phi(x,\mu)\\ &+\tilde{\mathbb{E}}\left[\frac{1}{2}\partial_{\tilde{x}}\partial_{\mu}\Phi(x,\mu,\tilde{\xi})+\partial_{\mu}\Phi(x,\mu,\tilde{\xi})\partial_{y}H(\tilde{\xi},\mu,\Psi(\tilde{\xi},\mu))\right],\end{split}$ (6.37)
@@
A function $\Psi\in C^{2,1}(\mathbb{R}\times\mathcal{P}_{2})$ is said to be a test function if the quantities:
[…]
and
[…]
are finite, uniformly in $(x,\mu)$ in any compact subset of $\mathbb{R}\times\mathcal{P}_{2}$.
Definition 6.7
Let $\mathcal{U}\in C(\mathbb{R}\times\mathcal{P}_{2})$. Then $\mathcal{U}$ is called a viscosity subsolution (resp. supersolution) of PDE (6.36) if, whenever $\Psi$ is a test function, and $(x^{0},\mu^{0})\in\mathbb{R}\times\mathcal{P}_{2}$ is a local maximum (resp. minimum) of $\mathcal{U}-\Psi$, we have
$(\mathcal{L}\Psi)[{\cal U}](x^{0},\mu^{0})+F(x^{0},\mu^{0},\mathcal{U}(x^{0},\mu^{0}))\geq 0,$ (6.41)
@@
$(\mathcal{L}\Psi)[{\cal U}](x^{0},\mu^{0})+F(x^{0},\mu^{0},\mathcal{U}(x^{0},\mu^{0}))\leq 0,$ (6.42)
The function $\mathcal{U}$ is called a viscosity solution of PDE (6.36) if it is both a viscosity subsolution and a viscosity supersolution.
@@
Due to the uniqueness for solutions of FBSDE, it is not hard to see that for any $t\geq 0$, $\mathcal{V}(X_{t}^{x,\xi},\mathcal{L}_{X_{t}^{\xi}})=Y_{t}^{x,\xi}$. Let $\Psi$ be a test function, and $(x^{0},\mu^{0})\in\mathbb{R}\times\mathcal{P}_{2}$ be a local maximum of $\mathcal{U}-\Psi$. We assume without loss of generality that $\mathcal{V}(x^{0},\mu^{0})=\Psi(x^{0},\mu^{0})$. We suppose that
$(\mathcal{L}\Psi)[\mathcal{V}](x^{0},\mu^{0})+F(x^{0},\mu^{0},\mathcal{V}(x^{0},\mu^{0}))<0,$ (6.43)
@@
$\begin{cases}\mathcal{V}(x,\mu)\leq\Psi(x,\mu),\\ (\mathcal{L}\Psi)[\mathcal{V}](x,\mu)+F(x,\mu,\mathcal{V}(x,\mu))<0.\end{cases}$ (6.44)
Taking a initial state $\xi^{0}\in\mathbb{L}^{2}(\mathcal{F}_{0};\mu^{0})$, we consider the processes $(X_{t}^{\xi^{0}},Y_{t}^{\xi^{0}},Z_{t}^{\xi^{0}})$ and $(X_{t}^{x^{0},\xi^{0}},Y_{t}^{x^{0},\xi^{0}},Z_{t}^{x^{0},\xi^{0}})$ which are solutions to FBSDEs (6.1) and (6.2). For some $T>0$, let $\tau$ denote the stopping time
$\tau\triangleq\inf\{t>0|(X_{t}^{x^{0},\xi^{0}},\mathcal{L}_{X_{t}^{\xi^{0}}})\notin O\}\land T.$ (6.45)
Let $\rho_{t}\triangleq{\cal L}_{X_{t}^{\xi^{0}}}$, we first note that the pair of processes
@@
$\begin{split}\bar{Y}_{t}=&\mathcal{V}(X^{x^{0},\xi^{0}}_{\tau},\rho_{\tau})+\int_{t}^{T}I_{[0,\tau]}(s)F(X^{x^{0},\xi^{0}}_{s},\rho_{s},\mathcal{V}(X^{x^{0},\xi^{0}}_{s},\rho_{s}))\mathrm{d}s\\ &-\int_{t}^{T}\bar{Z}_{s}\mathrm{d}B_{s},\quad\quad 0\leq t\leq T.\end{split}$ (6.47)
@@
$(\hat{Y}_{t},\hat{Z}_{t})\triangleq(\Psi(X^{x^{0},\xi^{0}}_{t\wedge\tau},\rho_{t\wedge\tau}),I_{\left[0,\tau\right]}(t)\partial_{x}\Psi(X^{x^{0},\xi^{0}}_{t\wedge\tau},\rho_{t\wedge\tau}))$ (6.48)
@@
$\begin{split}\hat{Y}_{t}=&\Psi(X^{x^{0},\xi^{0}}_{\tau},\rho_{\tau})-\int_{t}^{T}I_{[0,\tau]}(s)\mathcal{L}\Psi[\mathcal{V}](X^{x^{0},\xi^{0}}_{s},\rho_{s})\mathrm{d}s\\ &-\int_{t}^{T}\hat{Z}_{s}\mathrm{d}B_{s},\quad\quad 0\leq t\leq T.\end{split}$ (6.49)
@@
$\beta_{s}=-\mathcal{L}\Psi[\mathcal{V}](X^{x^{0},\xi^{0}}_{s},\rho_{s})-F(X^{x^{0},\xi^{0}}_{s},\rho_{s},\mathcal{V}(X^{x^{0},\xi^{0}}_{s},\rho_{s}))$ (6.50)
@@
$\begin{split}\tilde{Y}_{t}=&\Psi(X^{x^{0},\xi^{0}}_{\tau},\rho_{\tau})-\mathcal{V}(X^{x^{0},\xi^{0}}_{\tau},\rho_{\tau})+\int_{t}^{T}I_{[0,\tau]}(s)\beta_{s}\mathrm{d}s\\ &-\int_{t}^{T}\tilde{Z}_{s}\mathrm{d}B_{s},\quad\quad 0\leq t\leq T.\end{split}$ (6.52)
@@
$\tilde{Y}_{\tau}=\Psi(X^{x^{0},\xi^{0}}_{\tau},\rho_{\tau})-\mathcal{V}(X^{x^{0},\xi^{0}}_{\tau},\rho_{\tau})\geq 0,\quad\beta_{s}>0,\quad s\in[0,\tau].$ (6.54)